a) You need to make an aqueous solution of 0.194 M sodium phosphate for an experiment in lab, using a 250 mL volumetric flask. How much solid sodium phosphate should you add?

b)How many milliliters of an aqueous solution of 0.229 M ammonium iodide is needed to obtain 6.59 grams of the salt?
Answer #1

a)

volume , V = 250 mL

= 0.250 L

use:

number of mol,

n = Molarity * Volume

= 0.194*0.25

= 4.85*10^-2 mol

Molar mass of Na3PO4,

MM = 3*MM(Na) + 1*MM(P) + 4*MM(O)

= 3*22.99 + 1*30.97 + 4*16.0

= 163.94 g/mol

use:

mass of Na3PO4,

m = number of mol * molar mass

= 4.85*10^-2 mol * 1.639*10^2 g/mol

= 7.951 g

Answer: 7.95 g

b)

Molar mass of NH4I,

MM = 1*MM(N) + 4*MM(H) + 1*MM(I)

= 1*14.01 + 4*1.008 + 1*126.9

= 144.942 g/mol

mass(NH4I)= 6.59 g

use:

number of mol of NH4I,

n = mass of NH4I/molar mass of NH4I

=(6.59 g)/(1.449*10^2 g/mol)

= 4.547*10^-2 mol

use:

M = number of mol / volume in L

0.229 = 4.547*10^-2/ volume in L

volume = 0.199 L

volume = 199 mL

Answer: 199 mL